solve by the substition method
7x+3y=-35
-7x+y=63
the soltion is
or there is no solution
--------------------------------------------------------------------------------

Answer by edjones(8007)

7x+3y=-35
-7x+y=63
----------add
4y=28
.
y=7
x=-8
.
Ed